2. If the price of°ffice visits increases by 20% and the number of visits per family per year declines by 10%, what is the price elasticity of demand? Is consumption for office visits price elastic, inelastic, or unit elastic? (2 points)
Answer #1

Original price of office visit = P

New price of office visit = P (1+20%) = 1.2P

Original number of visits = Q

New number of visits = Q(1-10%) = 0.9Q

Average number of visits = (0.9Q+Q)/2 = 0.95Q

Average price = (1.2P+P)/2 = 1.1P

% Change in number of visits = ((0.9Q-Q)/0.95Q )*100 = -10.52%

% change in price = ((1.2P-P)/1.1P)*100 = 18.18%

Price elasticity of demand = % change in number of visits/% change in price

Price elasticity of demand = -10.52/18.18 = -0.58

As the price elasticity of demand is less than 1 which means it is inelastic.